hhanan:

What is the general equation (variables only) of the axis of symmetry for a quadratic graph.

Downpour17:

Do you still need help?

hhanan:

yes

Downpour17:

I can't help you, but this guy can @Convert

jhonyy9:

like a first step do you know what is the graph of a quadratic ?

hhanan:

no

Convert:

In standard form the general equation for an axis of symmetry is y=ax2+bx+c

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert In standard form the general equation for an axis of symmetry is y=ax2+bx+c \(\color{#0cbb34}{\text{End of Quote}}\) good job

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert In standard form the general equation for an axis of symmetry is y=ax2+bx+c \(\color{#0cbb34}{\text{End of Quote}}\) good job \(\color{#0cbb34}{\text{End of Quote}}\) Ty jhonyy (:

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert In standard form the general equation for an axis of symmetry is y=ax2+bx+c \(\color{#0cbb34}{\text{End of Quote}}\) good job \(\color{#0cbb34}{\text{End of Quote}}\) Ty jhonyy (: \(\color{#0cbb34}{\text{End of Quote}}\) np yw